OpenStudy (anonymous):

25.0mL of 0.1M Acetic acid is titrated to the equivalence point with 25.0mL of 0.1M NaOH. Calculate the pH at the equivalence point. Ka for acetic acid is 1.8 x 10^-5

OpenStudy (anonymous):

I.) Since you have have equal concentration of both acetic acid and NaOH .025L, 0.1M for both, just use kb and [H], and ph 1.)ka= 1.8*10^-5 kb= 1.0*10^-14 1.8^-5 =5.67*10^-6 2.) [H]= 1.0*10^-14 sqrt5.67*10^-6*.025 =1.92*10^-9 3.) log(1.92*10^-9) pH=8.72